OpenStudy (anonymous):

Find\[\sum_{k=n+1}^{3n}(3k+2)\]

OpenStudy (anonymous):

start from 1 to 3n then subtract what you get from 1 to n

OpenStudy (anonymous):

\[\sum_{k=1}^{3n}-\sum_{k=1}^n\]

OpenStudy (experimentx):

\[ 3 \sum_{k=n+1}^{3n}k + 2\sum_{k=n+1}^{3n} 1\\ 3( {3n+1 - (n+1)\over 2} ) (n+1 + 3n) + 2 (3n+1 - (n+1))\]

OpenStudy (anonymous):

Is it 12n^2+4n

OpenStudy (anonymous):

Total number of terms is 2n

OpenStudy (anonymous):

after algebra, i got \(12n^2+7n\) maybe i messed up is there a slicker way to do it?

hartnn (hartnn):

maybe substitute k-n = p ....

hartnn (hartnn):

what do u suggest @siddhantsharan ?

OpenStudy (anonymous):

Are you guys sure about the answer?

OpenStudy (anonymous):

I think its either 12n^2 +7n or 12n^2+4n

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

satellite73 is correct

OpenStudy (anonymous):